The vapor pressure of liquid C4H10 is 324.1 mmHg at a temperature of 251.0 K. The enthalpy of vaporization, ΔHvap, for this liquid is 22.4 kJ/mol.

What is the vapor pressure of C4H10 at a temperature of 269.2 K, in mmHg?
Answer #1

Apply Clasius Clapeyron equations

ln(P2/P1) = H/R*(1/T1 - 1/T2)

ln(P2/324.1) = 22400/8.313*(1/251 - 1/269.2)

Solve for P2

ln(P2/324.1) = 0.725793

P2 = exp(0.725793)*324.1 = 669.7102 mm Hg